# SkewRoute: Training-Free LLM Routing for Knowledge Graph Retrieval- Augmented Generation via Score Skewness of Retrieved Context

### v1

@@ 4.2 Results
Figure 5 present the experimental results of our four routing methods on the WebQSP dataset. Overall, all skewness indicators in our training-free framework consistently outperform the random mixing baseline across all scenarios on Hit@1 and F1-Score. Although the naive area-based method is less effective than other methods, its performance continuously remains above the baseline. Specifically, when routing between Qwen7b and Qwen72b, our methods cut larger LLM call ratio by half while achieving Hit@1 and F1-Score on par with full inference of larger LLM. Beyond this point, they even surpass larger LLM-only performance. Furthermore, we also test with Llama8b and Llama70b. Exemplified by the cumulative threshold-based routing method, a performance gain of nearly 2% in both Hit@1 and F1-score can be observed compared to random mixing at larger LLM call ratio nearly 0.2. These results strongly demonstrate the effectiveness of our routing methods based on score skewness to optimizing performance and cost. Additionally, we observe that the performance gap between Qwens is smaller than that between Llamas. While routing is effective in both cases, the narrower gap produces superior routing outcomes.
Figure 6 illustrates our experimental results on the CWQ, which is a more challenging KGQA dataset. Overall, our training-free methods possess and strong flexibility and generalization, performing well across various evaluation metrics and model combinations in new data scenario. Gini coefficient-based, entropy-based, and cumulative threshold-based routing methods deliver comparable performance and outperform the naive area-based method, which suffers from scaling sensitivity and capturing distribution feature. For instance, in Figure 6(c), within the larger LLM call ratio range from 0.3 to 0.6, the area-based method already exceeds the baseline, while the other methods achieve 1% higher performance than it. This underscores the importance of effectively capturing distribution characteristics.
Figure 7: Results of Routing Between Multiple Models on CWQ Dataset.
(a) WebQSP Hit@1
(b) WebQSP F1-Score
(c) CWQ Hit@1
(d) CWQ F1-Score
Figure 8: Results of Routing Between Qwen7b and Llama70b. (a)(b) on WebQSP and (c)(d) on CWQ.
4.3 Analysis
@@ 4.3.1 Routing Across Diverse Model Sizes
In this section, we further extend our methods to include an additional model size to demonstrate its generalization. Specifically, we use Qwen7b as the small LLM, Qwen72b as the large LLM, and Qwen14b as the medium LLM, which sits between them in terms of parameters, performance, and cost. The experimental results are presented in Figure 7. Notably, including the medium LLM significantly improves the trade-off between performance and cost. Routing with the medium model yields much better results than routing directly between the small and large models. Overall, our method also demonstrates clear effectiveness, consistently outperforming the random mixing baseline. This highlights the generalization and robustness of our approach, effectively enabling training-free routing across multiple LLM sizes and allowing for finer-grained cost-performance trade-offs tailored to the user’s budget.
@@ 4.3.2 Routing Across Model Families
As shown in Figure 8, we further explore routing across different model families, such as between Qwen7b and Llama70b, to evaluate the generalization ability of our method. Even when routing across distinct model families, our routing framework remains consistently effective, demonstrating strong robustness and generalization. Specifically, in terms of Hit@1 and F1-Score on the WebQSP dataset, our method achieves nearly a 3% improvement compared to the random-mixing baseline, while maintaining a cost increase of only 5%.
4.3.3 Routing Across Cumulative Probability
(a) WebQSP Hit@1
(b) WebQSP F1-Score
(c) CWQ Hit@1
(d) CWQ F1-Score
@@ 6 Conclusion
In this paper, we present the first LLM routing framework tailored for KG-RAG in a plug-and-play manner. This is motivated by the observation that the skewness of retrieval score distributions naturally reflects query difficulty in reasoning over external knowledge, aligning well with the RAG paradigm. By exploiting this skewness, our method dynamically routes queries among LLMs of different scales, achieving an effective trade-off between performance and cost. This simple yet robust solution also overcomes prior methods’ dependence on costly training data and generalizes well across datasets and model sizes. Extensive experiments validate the effectiveness of our approach. Our method offers a new perspective for efficient RAG deployment by reducing costs through LLM routing. Future work will focus on extending this training-free routing paradigm to broader RAG settings, including text-based and multimodal scenarios.
@@ A.1 Datasets
Table 1 summarizes the statistics of the datasets used in our experiments. For simplicity, we define one knowledge context critical for reasoning as a single hop. Specifically, the WebQSP dataset consists of 2,826 training samples and 1,628 test samples, with a maximum hop count of 2. The CWQ dataset contains 27,639 training samples and 3,531 test samples, with a maximum hop count of 4.
@@
A.2 Performance and Cost
@@
Driven by architectural innovations and scaling laws, LLMs has emerged. As shown in the Table 3 and Table 3, the cost-efficient smaller LLMs have limited capabilities while powerful larger LLMs incur high inference cost. This inherent performance-cost trade-off poses a significant deployment challenge. LLM routing offers an effective solution by directing queries to balance response quality with cost efficiency. For users, routing enables them handling simpler queries on their devices locally while call to expensive cloud APIs only for more difficult queries. For LLM platform providers, routing facilitates automatically directing simple queries to lower-cost smaller LLMs on the backend without sacrificing user experience. Consequently, LLM routing provides a flexible and cost-effective method, helping to unlock the full potential of LLMs across various scales while accommodating diverse budget and performance requirements.
A.3 Analysis Experiments
A.3.1 Score Skewness
Figure 10 illustrates the scores of retrieved contexts for queries within the WebQSP dataset, a trend likewise observed in the CWQ dataset. Specifically, the left panel displays high skewness, with retrieved context scores declining sharply as rank increases. This highly skewed scores indicate that a few of high-ranked contexts contribute the vast majority of the total score, while the extensive long tail comprises numerous contexts with significantly lower individual scores. In stark contrast, the right panel shows low skewness, deviating from this linear drop pattern. A prominent inflection point in its upper-right region indicates that contexts, particularly those of higher scores, diminish at a substantially more gradual pace than that of high skewness.
@@ A.3.1 Score Skewness
A.3.2 Cumulative Frequency Distribution
Building upon the previously observed variations in score skewness across queries, we visualized their cumulative frequency distributions on the WebQSP and CWQ dataset, as shown in Figure 11. These plots reveal two primary types of cumulative behavior: high skewness and low skewness. For the former, where scores are concentrated among just a few retrieved contexts, the cumulative frequency rises steeply. Thus, a high cumulative probability is achieved with only a minimal number of contexts (e.g., Figure 11(a) illustrates reaching 95% cumulative probability with just 5 triples). In contrast, the low skewness where scores are more evenly distributed among more contexts, exhibit a slower rise in cumulative frequency, requiring more to reach the same probability (e.g., Figure 11(b), 95% requires 63). Similar behavior was also observed in the CWQ dataset. This significant variation in the number of contexts required for a given cumulative probability directly reflects the differences in underlying distribution skewness of scores. This observation inspires our cumulative threshold-based method, which leverages the context count needed to achieve a preset cumulative probability to inform dynamic routing decisions.

### v2

@@ 4.2 Results
Dataset Method 0% 20% 40% 60% 80% 100% Avg. Eff.
WebQSP Random Routing 77.52 78.18 78.85 79.51 80.18 80.84 -
RouteLLM [ICLR, 25] 77.52 78.56 (+0.38) 79.18 (+0.33) 79.55 (+0.04) 79.98 (-0.20) 80.84 +0.14
GraphRouter [ICLR, 25] 77.52 78.26 (+0.08) 79.30 (+0.45) 80.22 (+0.71) 80.71 (+0.53) 80.84 +0.44
Ours (Gini-based) 77.52 79.48 (+1.30) 79.98 (+1.13) 81.20 (+1.69) 80.96 (+0.78) 80.84 +1.23
Ours (Entropy-based) 77.52 79.18 (+1.00) 80.34 (+1.49) 81.08 (+1.57) 80.77 (+0.59) 80.84 +1.16
Ours (Cumulative-based) 77.52 79.24 (+1.06) 80.04 (+1.19) 80.71 (+1.20) 80.71 (+0.53) 80.84 +1.00
CWQ Random Routing 45.68 47.59 49.51 51.42 53.34 55.25 -
RouteLLM [ICLR, 25] 45.68 47.81 (+0.22) 50.41 (+0.90) 52.00 (+0.58) 53.61 (+0.27) 55.25 +0.49
GraphRouter [ICLR, 25] 45.68 47.83 (+0.24) 50.10 (+0.59) 52.02 (+0.60) 53.78 (+0.44) 55.25 +0.47
Ours (Gini-based) 45.68 48.94 (+1.35) 50.92 (+1.41) 52.53 (+1.11) 54.15 (+0.81) 55.25 +1.17
Ours (Entropy-based) 45.68 48.74 (+1.15) 50.72 (+1.21) 52.70 (+1.28) 54.01 (+0.67) 55.25 +1.08
Ours (Cumulative-based) 45.68 49.02 (+1.43) 50.89 (+1.38) 52.82 (+1.40) 53.89 (+0.55) 55.25 +1.19
Table 1: Routing Between Qwen2.5 Models: 7B as the Small LLM and 72B as the Large LLM.
@@
Table 1 present the experimental results of our routing methods on the WebQSP and CWQ datasets between Qwen7B and Qwen72B. Overall, all skewness indicators in our training-free framework consistently outperform baselines across all scenarios. Specifically, our gini coefficient-based, entropy-based and cumulative threshold-based methods cut larger LLM call ratio by 40% while achieving on par with full inference of larger LLM. Furthermore, the gini coefficient-based routing method demonstrates a significant average effectiveness, outperforming RouteLLM by over 8x and surpassing another strong baseline, GraphRouter, by more than 2x on WebQSP dataset. The same trend is observed on the CWQ dataset, where our method shows consistent improvements over the baselines. These results strongly demonstrate the effectiveness of our routing methods based on score skewness to optimizing performance and cost. Additionally, we find that the performance gap on WebQSP is smaller than that of CWQ dataset. While routing is effective in both cases, the narrower gap produces superior routing outcomes.
Dataset Method 0% 20% 40% 60% 80% 100% Avg. Eff.
WebQSP Random Routing 78.56 79.68 80.80 81.91 83.03 84.15 -
RouteLLM [ICLR, 25] 78.56 79.73 (+0.05) 80.47 (-0.33) 81.82 (-0.09) 82.86 (-0.17) 84.15 -0.14
GraphRouter [ICLR, 25] 78.56 79.67 (-0.01) 80.65 (-0.15) 81.39 (-0.52) 83.35 (+0.32) 84.15 -0.09
Ours (Gini-based) 78.56 81.33 (+1.65) 81.57 (+0.77) 82.62 (+0.71) 83.35 (+0.32) 84.15 +0.86
Ours (Entropy-based) 78.56 81.08 (+1.40) 82.00 (+1.20) 82.49 (+0.58) 83.66 (+0.63) 84.15 +0.95
Ours (Cumulative-based) 78.56 81.57 (+1.89) 81.27 (+0.47) 82.62 (+0.71) 83.60 (+0.57) 84.15 +0.91
CWQ Random Routing 49.90 51.51 53.12 54.72 56.33 57.94 -
RouteLLM [ICLR, 25] 49.90 51.26 (-0.25) 53.61 (+0.49) 55.03 (+0.31) 56.75 (+0.42) 57.94 +0.24
GraphRouter [ICLR, 25] 49.90 50.84 (-0.67) 52.79 (-0.33) 54.80 (+0.08) 56.16 (-0.17) 57.94 -0.27
Ours (Gini-based) 49.90 52.65 (+1.14) 55.00 (+1.88) 56.16 (+1.44) 57.04 (+0.71) 57.94 +1.29
Ours (Entropy-based) 49.90 52.51 (+1.00) 54.89 (+1.77) 56.07 (+1.35) 56.78 (+0.45) 57.94 +1.14
Ours (Cumulative-based) 49.90 52.68 (+1.17) 54.77 (+1.65) 56.41 (+1.69) 56.61 (+0.28) 57.94 +1.20
Table 2: Routing Between Llama3.1 Models: 8B as the Small LLM and 70B as the Large LLM.
Dataset Method 0% 20% 40% 60% 80% 100% Avg. Eff.
WebQSP Random Routing 77.52 78.85 80.17 81.50 82.82 84.15 -
RouteLLM [ICLR, 25] 77.52 78.81 (-0.04) 79.79 (-0.38) 81.33 (-0.17) 82.49 (-0.33) 84.15 -0.23
GraphRouter [ICLR, 25] 77.52 78.99 (+0.14) 80.84 (+0.67) 82.06 (+0.56) 83.42 (+0.60) 84.15 +0.49
Ours (Gini-based) 77.52 81.51 (+2.66) 82.49 (+2.32) 83.66 (+2.16) 83.97 (+1.15) 84.15 +2.07
Ours (Entropy-based) 77.52 81.27 (+2.42) 82.74 (+2.57) 83.42 (+1.92) 83.97 (+1.15) 84.15 +2.02
Ours (Cumulative-based) 77.52 81.63 (+2.78) 82.31 (+2.14) 83.60 (+2.10) 83.85 (+1.03) 84.15 +2.02
CWQ Random Routing 45.68 48.13 50.58 53.04 55.49 57.94 -
RouteLLM [ICLR, 25] 45.68 48.12 (-0.01) 51.09 (+0.51) 53.58 (+0.54) 56.05 (+0.56) 57.94 +0.40
GraphRouter [ICLR, 25] 45.68 48.00 (-0.13) 50.78 (+0.20) 53.50 (+0.46) 55.31 (-0.18) 57.94 +0.09
Ours (Gini-based) 45.68 50.13 (+2.00) 53.04 (+2.46) 54.83 (+1.79) 56.53 (+1.04) 57.94 +1.82
Ours (Entropy-based) 45.68 49.82 (+1.69) 52.82 (+2.24) 54.86 (+1.82) 56.41 (+0.92) 57.94 +1.67
Ours (Cumulative-based) 45.68 49.90 (+1.77) 52.76 (+2.18) 55.00 (+1.96) 56.05 (+0.56) 57.94 +1.62
Table 3: Routing Across Model Families: Qwen2.5-7B-Instruct and Llama-3.1-70B-Instruct.
Table 2 illustrates our experimental results routing between Llama8B and Llama70B. Overall, our training-free methods possess and strong flexibility and generalization, performing well across various datasets and model combinations. Gini coefficient-based, entropy-based, and cumulative threshold-based routing methods demonstrate a significant advantage. Even under conditions where RouteLLM and GraphRouter fail completely, with their performance dropping below the naive random routing, our routing methods deliver a notable gain, increasing the average effectiveness by nearly 1% on WebQSP dataset. On the more challenging CWQ dataset, our routing methods overwhelmingly surpass all baselines. While the GraphRouter still remains ineffective, our gini coefficient-based routing method achieves average effectiveness 5x better than RouteLLM.
Dataset Method 0% 20% 40% 60% 80% 100% Avg. Eff.
WebQSP Random Routing 72.79 73.74 74.68 75.63 76.57 77.52 -
RouteLLM [ICLR, 25] 72.79 74.26 (+0.52) 75.25 (+0.57) 76.17 (+0.54) 77.33 (+0.76) 77.52 +0.60
GraphRouter [ICLR, 25] 72.79 73.59 (-0.15) 74.69 (+0.01) 75.92 (+0.29) 77.03 (+0.46) 77.52 +0.15
Ours (Gini-based) 72.79 75.12 (+1.38) 77.27 (+2.59) 76.78 (+1.15) 77.52 (+0.95) 77.52 +1.52
Ours (Entropy-based) 72.79 75.31 (+1.57) 77.21 (+2.53) 77.21 (+1.58) 77.46 (+0.89) 77.52 +1.64
Ours (Cumulative-based) 72.79 75.25 (+1.51) 77.21 (+2.53) 77.58 (+1.95) 77.40 (+0.83) 77.52 +1.71
CWQ Random Routing 41.18 42.08 42.98 43.88 44.78 45.68 -
RouteLLM [ICLR, 25] 41.18 41.97 (-0.11) 42.65 (-0.33) 43.39 (-0.49) 44.41 (-0.37) 45.68 -0.33
GraphRouter [ICLR, 25] 41.18 41.49 (-0.59) 42.03 (-0.95) 43.16 (-0.72) 44.72 (-0.06) 45.68 -0.58
Ours (Gini-based) 41.18 42.45 (+0.37) 43.59 (+0.61) 44.94 (+1.06) 45.37 (+0.59) 45.68 +0.66
Ours (Entropy-based) 41.18 42.65 (+0.57) 43.64 (+0.66) 44.94 (+1.06) 45.20 (+0.42) 45.68 +0.68
Ours (Cumulative-based) 41.18 42.31 (+0.23) 43.50 (+0.52) 44.86 (+0.98) 45.51 (+0.73) 45.68 +0.62
Table 4: Routing of Retrieved Contexts Nums (Top-K, k=10 vs.k=100) of Qwen2.5-7B-Instruct.
4.3 Case Study
@@ 4.3.1 Routing Across Diverse Model Sizes
(a) WebQSP Dataset.
(b) CWQ Dataset.
Figure 5: Routing Between Multiple Models
In this section, we further extend our methods to include an additional model size to demonstrate its generalization. Specifically, we use Qwen7B as the small LLM, Qwen72B as the large LLM, and Qwen14B as the medium LLM, which sits between them in terms of parameters, performance, and cost. The experimental results are presented in Figure 5. Notably, including the medium LLM significantly improves the trade-off between performance and cost. Routing with the medium model yields much better results than routing directly between the small and large models. Overall, our method also demonstrates clear effectiveness, consistently outperforming the random routing baseline. This highlights the generalization and robustness of our methods, effectively enabling training-free routing across multiple LLM sizes and allowing for finer-grained cost-performance trade-offs tailored to the user’s budget.
@@ 4.3.2 Routing Across Model Families
As shown in Table 3, we further explore routing across different model families, such as between Qwen7B and Llama70B, which are trained on different corpora and have different model architectures, to evaluate the generalization ability of our routing methods. Even when routing across distinct model families, our routing framework remains consistently effective, demonstrating strong robustness and generalization. Specifically, on the WebQSP dataset, the average effectiveness of the gini coefficient-based routing method is over 4x higher than GraphRouter. On the CWQ dataset, our routing methods significantly outperform all baselines. At a 40% large LLM call ratio, our gini coefficient-based routing method brings a 2.57% and 2.46% improvement over random routing on the WebQSP and CWQ datasets, respectively. These results highlight the superiority of our routing methods in reducing inference costs while maintaining inference performance. Our gini coefficient-based routing method achieves the best average effectiveness. Therefore, we recommend it for routing across model families.
4.3.3 Routing of Retrieved Contexts Nums
In addition to routing between models of different sizes, dynamically adjusting the number of retrieved contexts is another promising direction. For simple queries, reducing the number of retrieved contexts effectively decreases the input token count and lowers inference costs. In this case study, we employ our skewness-based routing algorithm to achieve this: for complex queries, we retain 100 retrieved contexts as before, whereas for simple queries, only 10 contexts are used. As illustrated in Table 4, our routing methods demonstrate remarkable efficiency and performance. Notably, even when using only 10 retrieved contexts for 60% of the queries, our cumulative threshold-based routing method outperformes using 100 contexts for all queries on WebQSP, while drastically reducing inference costs. Furthermore, the average effectiveness of this method is nearly 3x that of RouteLLM. This highlights the broad applicability of our proposed skewness-based plug-and-play algorithm and its promising potential for future applications.
4.3.4 Other Scorer in KG-RAG
The focus of this work is not on improving the scorers used in the KG-RAG pipeline. Instead, our method leverages the score distributions produced by existing scorers in a plug-and-play manner for routing, making it largely agnostic to the choice of scorer. To further demonstrate the generality of our method, we conduct additional evaluations using a domain fine-tuned MLP scorer in conjunction with the gte-large-en-v1.5 embedding model as new scorer. Across all settings, our routing method still achieves substantial performance gains over existing routing methods. These results provide further validation of the effectiveness and broad applicability of our methods, as shown in the Table 5. Although the scorer trained on the semantic similarity between context and query don’t perform as well as SubgraphRAG, our routing methods still lead the pack, a strong testament to their versatility. Specifically, our entropy-based routing method is particularly notable, with its average effectiveness being nearly 3x that of RouteLLM.
Dataset Method 0% 20% 40% 60% 80% 100% Avg. Eff.
WebQSP Random Routing 75.61 76.42 77.23 78.05 78.86 79.67 -
RouteLLM [ICLR, 25] 75.61 76.66 (+0.24) 77.40 (+0.17) 78.01 (-0.04) 78.50 (-0.36) 79.67 +0.003
GraphRouter [ICLR, 25] 75.61 76.84 (+0.42) 77.95 (+0.72) 78.93 (+0.88) 80.04 (+1.18) 79.67 +0.80
Ours (Gini-based) 75.61 77.46 (+1.04) 78.62 (+1.39) 79.55 (+1.50) 79.67 (+0.81) 79.67 +1.19
Ours (Entropy-based) 75.61 77.40 (+0.98) 78.69 (+1.46) 79.55 (+1.50) 79.73 (+0.87) 79.67 +1.20
Ours (Cumulative-based) 75.61 77.58 (+1.16) 78.56 (+1.33) 79.79 (+1.74) 79.61 (+0.75) 79.67 +1.25
CWQ Random Routing 43.98 45.72 47.46 49.20 50.94 52.68 -
RouteLLM [ICLR, 25] 43.98 46.13 (+0.41) 48.12 (+0.66) 49.70 (+0.50) 51.29 (+0.35) 52.68 +0.48
GraphRouter [ICLR, 25] 43.98 45.71 (-0.01) 47.78 (+0.32) 49.67 (+0.47) 51.29 (+0.35) 52.68 +0.28
Ours (Gini-based) 43.98 46.84 (+1.12) 49.05 (+1.59) 50.89 (+1.69) 51.97 (+1.03) 52.68 +1.36
Ours (Entropy-based) 43.98 46.87 (+1.15) 49.16 (+1.70) 51.06 (+1.86) 51.85 (+0.91) 52.68 +1.41
Ours (Cumulative-based) 43.98 46.45 (+0.73) 48.99 (+1.53) 51.01 (+1.81) 51.91 (+0.97) 52.58 +1.26
Table 5: Routing Between Qwen Models with Other Scorer: 7B as the Small LLM and 72B as the Large LLM.
4.3.5 Computational Efficiency of Routing
Table 6 presents a runtime comparison of different routing methods. Our method is training-free and extremely fast, as it estimates the skewness level simply by analyzing the score distribution of the top-$K$ retrieved contexts. In contrast, training-based methods like GraphRouter and RouteLLM require GPU resources for deployment and suffer significant slowdowns when only CPUs are available. Our method, however, runs efficiently on CPUs alone. Specifically, GraphRouter incurs substantial latency due to query embedding, running 5,460× slower than our cumulative-based method on CPUs, and still 2,714× slower even with GPU acceleration.
Model RouteLLM GraphRouter Our (Gini-based) Our (Entropy-based) Our (Cumulative-based) Our (Area-based)
CPU 25.21 46.41 0.01 0.03 0.01 0.01
GPU 10.12 23.07 - - - -
Table 6: Time Statistics of Different Routing Methods (ms/query).
@@ 6 Conclusion
In this paper, we present the first LLM routing framework tailored for KG-RAG in a plug-and-play manner. This is motivated by the observation that the skewness of retrieval score distributions naturally reflects query difficulty in reasoning over external knowledge, aligning well with the RAG paradigm. By exploiting this skewness, our method dynamically routes queries among LLMs of different scales, achieving an effective trade-off between performance and cost. This simple yet robust solution also overcomes prior methods’ dependence on costly training data and generalizes well across datasets and model sizes. Extensive experiments validate the effectiveness of our method. We offer a new perspective for efficient RAG deployment by reducing costs through LLM routing. Future work will focus on extending this training-free routing paradigm to broader RAG settings, including text-based and multimodal scenarios.
@@ A.1 Datasets
Table 7 summarizes the statistics of the datasets used in our experiments. Specifically, the WebQSP dataset consists of 2,826 training samples and 1,628 test samples, with a maximum hop of 2. The CWQ dataset contains 27,639 training samples and 3,531 test samples, with a maximum hop count of 4.
@@
A.2 Implementation Details.
For reasoning, we use one-shot prompting to LLMs’ generation. Consistent with SubgraphRAG, the prompt template we employe is detailed in Appendix A.9. The temperature parameter is set to 0 for reproduction, and the maximum token length for generation is fixed at 4000.
A.3 Performance and Cost
@@
Driven by architectural innovations and scaling laws, LLMs has emerged. As shown in the Table 9, the cost-efficient smaller LLMs have limited capabilities while powerful larger LLMs incur high inference cost. This inherent performance-cost trade-off poses a significant deployment challenge. LLM routing offers an effective solution by directing queries to balance response quality with cost efficiency. For users, routing enables them handling simpler queries on their devices locally while call to expensive cloud APIs only for more difficult queries. For LLM platform providers, routing facilitates automatically directing simple queries to lower-cost smaller LLMs on the backend without sacrificing user experience. Consequently, LLM routing provides a flexible and cost-effective method, helping to unlock the full potential of LLMs across various scales while accommodating diverse budget and performance requirements.
A.4 Score Skewness on WebQSP Dataset
Figure 6 illustrates the scores of retrieved contexts for queries within the WebQSP dataset, a trend likewise observed in the CWQ dataset as described earlier. Specifically, the left panel displays high skewness, with retrieved context scores declining sharply as rank increases. This highly skewed scores indicate that a few of high-ranked contexts contribute the vast majority of the total score, while the extensive long tail comprises numerous contexts with significantly lower individual scores. In stark contrast, the right panel shows low skewness, deviating from this linear drop pattern. A prominent inflection point in its upper-right region indicates that contexts, particularly those of higher scores, diminish at a substantially more gradual pace than that of high skewness.
@@ A.4 Score Skewness on WebQSP Dataset
A.5 Cumulative Probability of Context Scores
Building upon the previously observed variations in score skewness across queries, we visualize their cumulative probability on the WebQSP and CWQ dataset, as shown in Figure 7. These plots reveal two primary types of cumulative behavior: high skewness and low skewness. For the former, where scores are concentrated among just a few retrieved contexts, the cumulative probability rises steeply. Thus, a high cumulative probability is achieved with only a minimal number of contexts (e.g., Figure 7(a) illustrates reaching 95% cumulative probability with just 5 triples). In contrast, the low skewness where scores are more evenly distributed among more contexts, exhibit a slower rise in cumulative probability, requiring more to reach the same value (e.g., Figure 7(b), 95% requires 63). Similar behavior was also observed in the CWQ dataset. This significant variation in the number of contexts required for a given cumulative probability directly reflects the differences in underlying distribution skewness of scores. This observation inspires our cumulative threshold-based method, which leverages the context number needed to achieve a preset cumulative probability to inform dynamic routing decisions.